OpenStudy (anonymous):

how do you get -46/7 from solving for x for (x-6)/(x+5)=8

OpenStudy (mayankdevnani):

\[\large \bf \frac{x-6}{x+5}=8\] Cross multiply it, \[\large \bf x-6=8x+40\]

OpenStudy (mayankdevnani):

and solve for x ? @E.St22

OpenStudy (anonymous):

thank you so much! @mayankdevnani

OpenStudy (mayankdevnani):

your welcome :)